Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2004-010_2004-016

Mission Operations Days: 2004/010 to 2004/016
January 10, 2004 00:00:00 GMT to January 16, 2004 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

AM-1 "Terra" is in Normal Mode
PFM MODIS is in A-Side Nominal Science Mode
(Cross Strapped with FR B)
ComponentState(s)Comment
BlackbodyA On (290K); B OffNominal Science Mode
Calibration ElectronicsA On; B OffNominal Science Mode
Control ProcessorA On; B OffNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ScreenedPermanent Configuration
FDDI FormatterA On; B OffNominal Science Mode
FDDI PortA On; B OffNominal Science Mode
FIFO Memory#1 & #2 On; #3 & #4 OffNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 On, #2 OffNominal Science Mode
PV VIS FPAA On; B OffNominal Science Mode
PV NIR FPAA On; B OffNominal Science Mode
PV SMIR FPAA On; B OffNominal Science Mode
PV LWIR FPAA On; B OffNominal Science Mode
PC LWIR FPAA On; B OffNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOn (83K)Nominal Science Mode
RC SMIR FPA HeaterOffNominal Science Mode
Scan AssemblyA On; B OffNominal Science Mode
SDSMOffNominal Science Mode
SRCAOffNominal Science Mode
PS1 Survival HeaterEnabledNominal Science Mode
PS2 Survival HeaterEnabledNominal Science Mode
Timing GeneratorA On; B OffNominal Science Mode
Flight SoftwareRev BD + 11 sets of patches
  • CP07 to Survival Mode - No SDD to closed command (1 patch)
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
  • CP07 to Safe Mode - No SDD to closed command (1 patch)
  • CP07 to Science Mode - Turn on FRB instead of FRA (1 patch)
  • CP07 to Outgas and Science Mode - Turn on FRB instead of FRA (1 patch)
  • FR Error Counting and Statistics (2 patches)
  • FR 1000ms time code patches (3 patches)
  • FR Band 5 Gain Change (20 commands)
  • FR BB Ideal Output Values (1 patch)
  • FR A EEPROM KME Fast 1 word patch (1 patch)
  • FR B EEPROM KME Fast 1 word patch (1 patch)
Nominal Science Mode
Inhibit Ids SetnoneNominal Science Mode
TMON 66DisabledNominal Science Mode
TMON 67DisabledNominal Science Mode
TMON 68EnabledNominal Science Mode
TMON 69DisabledNominal Science Mode

This Week’s Completed MODIS Activities:

Saturday, January 10, 2004

None

Sunday, January 11, 2004

None

Monday, January 12, 2004

Lunar Roll Maneuver (no roll required)
2004/012 07:11:10 ATC MOD_OA03_FR_DAY_RATE
2004/012 07:57:15 ATC MOD_OA03_FR_NIGHT_RATE (03:23 early)
2004/012 08:18:54-08:25:40 ATC MOD_OA08_ROLL_MAN
2004/012 08:53:26 ATC MOD_OA03_FR_DAY_RATE (03:23 late)
2004/012 09:39:31 ATC MOD_OA03_FR_NIGHT_RATE

Tuesday, January 13, 2004

2004/013 20:57:24-21:06:12 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement
2004/013 23:15:00 ATC MOD_SRCA_NOMINAL_SETUP.4 Sets up nominal SRCA lamp configuration
2004/013 23:18:07-00:16:18 ATC MOD_OA23_SRCA_FULL_SPAT_A.3 OA-23: SRCA Full Spatial Calibration - A Side
Data collection from 23:34:01-00:15:03

Wednesday, January 14, 2004

None

Thursday, January 15, 2004

None

Friday, January 16, 2004

None

This Week’s Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Upcoming January calibrations include weekly SDSM SCREENED, monthly SRCA 10W Radiometric, and quarterly SRCA Full Spectral.

Tuesday, January 20, 2004

2004/020 21:03:03-21:11:51 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement
2004/020 21:50:00 ATC MOD_SRCA_NOMINAL_SETUP.4 Sets up nominal SRCA lamp configuration
2004/020 21:52:46-22:13:41 ATC MOD_OA20_SRCA_10W_RAD_A_Base.2 OA-20: SRCA 10W Continuous Radiometric - A Side, Using Lamp #1
Data collection is from 22:08:40-22:13:41 (in S/C Night)

Maneuvers:

The date and time for the February Lunar Calibration has yet to be determined.

MODIS Anomalies:

MODIS NAD Anomaly:

On Wednesday, December 24, 2003 at 20:20:01 (DOY 2003/358), the MODIS Nadir Aperture Door (NAD) was commanded to OPEN. A situation that was expected from previous NAD movements, the "NAD Open" trigger (MOD_CR_DR_NAD_OPEN) did not register from the door movement (it did not change from "NOT_OPEN" to "OPEN" as one would expect). This was anticipated as a possibility. The "NAD Closed" trigger did perform as expected; several seconds after the CP noted it was moving the NAD, the trigger value changed from "CLOSED" to "NOT_CLOSED". This shows the door did move at least enough to allow that trigger to release. The PS1 input current profile looked nominal while the door driver was powered. The CP generated the expected xx69 "NAD Trigger Not Found" from the command door movement with lack of trigger response. However, the door does appear to be OPEN as seen from the data from the Rapid Response team. Science data also appears to be nominal.

The NAD OPEN trigger may yet “pop” to OPEN due to thermal cycling on the trigger, as was seen on 2002/079.

Terra SSR Anomaly:

Although not directly the result of a MODIS problem, the most recent Printed Wire Assembly (PWA) failures in the Terra Solid State Recorder (SSR) do affect MODIS data. See the Non-MODIS Significant Events section below for more details.

General Instrument Comments:

MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.

MODIS Telemetry Trends:

Nominal telemetry trending is ongoing.

Non-MODIS Significant Events:

Terra Solid State Recorder Anomaly

Overview:

The SSR Anomaly is still under investigation. MODIS currently is operating with 34 Supersets allotted for its buffer. There are plans in the works for possibly power cycling the SSR to get back PWA’s that are in the Bad List, but no date has yet been set. There are no current plans for a change in science data capture rates. For details, see previous reports.

Terra SFE Anomaly

Overview:

The Terra SFE-A was recycled at 2004/015 13:46 GMT. The window of lost science data during the recycle was 2004/015 13:43-13:53. A Direct Broadcast was observed at Goddard at 2004/015 14:21. The data looked normal. The recycle appears to have fixed the data corruption.

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.

http://jupiter02.gsfc.nasa.gov/links/links.htm

Limited Life Item Status:

SRCA 10W Lamp #1: 223.2 of 500 hours
SRCA 10W Lamp #2: 164.0 of 500 hours
SRCA 10W Lamp #3: 173.2 of 500 hours
SRCA 10W Lamp #4: 62.2 of 500 hours

SRCA 1W Lamp #1: 565.1 of 4000 hours
SRCA 1W Lamp #2: 276.7 of 4000 hours

Solar Diffuser Door: 2146 of 3022 Movements
Nadir Aperture Door: 538 of 1316 Movements
Space View Door: 441 of 1316 Movements